Find the eigenvalues and an eigenvector per eigenvalue of the matrix
A= [ 0 1 1]
[ 3 -4 -3]
[ -5 7 6]
Is A is diagonalisable ? Give reason your answer.
1
Expert's answer
2021-07-26T09:10:38-0400

"\text{Given the matrix \(A\) to find the Eigenvalues we have the following:}\\
{\rm det}(A - \lambda I) = 0\\
\implies \begin{vmatrix}- \lambda & 1 & 1\\3 & - \lambda - 4 & -3\\-5 & 7 & 6 - \lambda\end{vmatrix}\ = 0 \\
\text{Add column }2\text{ multiplied by λ to column 1:} C_{1} = C_{1} + \lambda C_{2}\\
\left|\begin{array}{ccc}- \lambda & 1 & 1\\3 & - \lambda - 4 & -3\\-5 & 7 & 6 - \lambda\end{array}\right| = \left|\begin{array}{ccc}0 & 1 & 1\\- \lambda \left(\lambda + 4\right) + 3 & - \lambda - 4 & -3\\7 \lambda - 5 & 7 & 6 - \lambda\end{array}\right| \\
\text{Subtract column }2\text{ from column 3: } C_{3} = C_{3} - C_{2} \\
\left|\begin{array}{ccc}0 & 1 & 1\\- \lambda \left(\lambda + 4\right) + 3 & - \lambda - 4 & -3\\7 \lambda - 5 & 7 & 6 - \lambda\end{array}\right| = \left|\begin{array}{ccc}0 & 1 & 0\\- \lambda \left(\lambda + 4\right) + 3 & - \lambda - 4 & \lambda + 1\\7 \lambda - 5 & 7 & - \lambda - 1\end{array}\right| \\
\text{Expand along row 1:} \\
\left|\begin{array}{ccc}0 & 1 & 0\\- \lambda \left(\lambda + 4\right) + 3 & - \lambda - 4 & \lambda + 1\\7 \lambda - 5 & 7 & - \lambda - 1\end{array}\right| = \left(0\right) \left(-1\right)^{1 + 1} \left|\begin{array}{cc}- \lambda - 4 & \lambda + 1\\7 & - \lambda - 1\end{array}\right| + \left(1\right) \left(-1\right)^{1 + 2} \left|\begin{array}{cc}- \lambda \left(\lambda + 4\right) + 3 & \lambda + 1\\7 \lambda - 5 & - \lambda - 1\end{array}\right| + \left(0\right) \left(-1\right)^{1 + 3} \left|\begin{array}{cc}- \lambda \left(\lambda + 4\right) + 3 & - \lambda - 4\\7 \lambda - 5 & 7\end{array}\right| = - \left|\begin{array}{cc}- \lambda \left(\lambda + 4\right) + 3 & \lambda + 1\\7 \lambda - 5 & - \lambda - 1\end{array}\right| \\
\text{Note that: } \\
\left|\begin{array}{cc}- \lambda \left(\lambda + 4\right) + 3 & \lambda + 1\\7 \lambda - 5 & - \lambda - 1\end{array}\right| = \left(- \lambda \left(\lambda + 4\right) + 3\right)\cdot \left(- \lambda - 1\right) - \left(\lambda + 1\right)\cdot \left(7 \lambda - 5\right) = \lambda^{3} - 2 \lambda^{2} - \lambda + 2 \\
\text{Finally we have the following: } \\
\left(-1\right)\cdot \left(\lambda^{3} - 2 \lambda^{2} - \lambda + 2\right) = - \left(\lambda - 2\right) \left(\lambda - 1\right) \left(\lambda + 1\right).\\
\implies {\rm det}(A - \lambda I) = - \left(\lambda - 2\right) \left(\lambda - 1\right) \left(\lambda + 1\right) =0 \\
\text{Solving this we have that: } \lambda = -1, 1,2.\\
\therefore \text{The eigenvalues of \(A\) are } \lambda = -1, \lambda = 1, \lambda = 2\\
\text{We now fin the eigenvectors of the matrix. }\\
1. \text{For } \lambda = 2:\\
\left[\begin{array}{ccc}- \lambda & 1 & 1\\3 & - \lambda - 4 & -3\\-5 & 7 & 6 - \lambda\end{array}\right] = \left[\begin{array}{ccc}-2 & 1 & 1\\3 & -6 & -3\\-5 & 7 & 4\end{array}\right]\\ 
\text{The nullspace of this matrix is } 
\left[\begin{array}{c}\frac{1}{3}\\- \frac{1}{3}\\1\end{array}\right] \text{which is the eigenvector.} \\
2. \text{For } \lambda = 1:\\
\left[\begin{array}{ccc}- \lambda & 1 & 1\\3 & - \lambda - 4 & -3\\-5 & 7 & 6 - \lambda\end{array}\right] = \left[\begin{array}{ccc}-1 & 1 & 1\\3 & -5 & -3\\-5 & 7 & 5\end{array}\right]\\
\text{The nullspace for this matrix is: } \left[\begin{array}{c}1\\0\\1\end{array}\right] \text{and this is the eigenvector} \\
3. \text{For }\lambda = -1: \\
\left[\begin{array}{ccc}- \lambda & 1 & 1\\3 & - \lambda - 4 & -3\\-5 & 7 & 6 - \lambda\end{array}\right] = \left[\begin{array}{ccc}1 & 1 & 1\\3 & -3 & -3\\-5 & 7 & 7\end{array}\right] \\
\text{The nullspace of this matrix is } \left[\begin{array}{c}0\\-1\\1\end{array}\right] \text{which is the eigenvector.}\\
\text{\(A\) is diagonalisable since each of the eigenvectors of \(A\) are linearly independent and they are distinct.}"